OpenStudy (anonymous):

it takes mandy 4 hrs to put up paneling in a room. omar takes 5 hrs to do the same job. how long would it take them, working together, to panel the room?

OpenStudy (anonymous):

in one hour mandy can do how much of the work???

OpenStudy (anonymous):

doesn't say.

hero (hero):

Hint: M = Mandy's Hours O = Omar's Hours \[Time = \frac{M \times O}{M + O}\]

OpenStudy (anonymous):

as he compltes the job in 4 hours so he will do 1/4th of the job in one hour

hero (hero):

\[Time = \frac{4 \times 5}{4 + 5}\]

OpenStudy (anonymous):

i see, thank you hero and sheg.

OpenStudy (anonymous):

similarly omar completes the job in 5 hours so he will complete 1/5th of the job in one hour

hero (hero):

By the way, when I say "Time", I mean in hours.

OpenStudy (anonymous):

got it, thanks.